OpenStudy (anonymous):

FAN+MEDAL If the sail ship is 10 feet long and has a diagonal length of 26 feet. What is the height of the sail in feet?

OpenStudy (gabebae):

Use the pythagorean theorem

OpenStudy (16shuston):

Do the same thing as last time. the Pythagorean theorem

OpenStudy (gabebae):

A^2 + B^2 = C^2

OpenStudy (16shuston):

it should look like this 10^2+B^2=26^2 solve for B

OpenStudy (coolguy34):

Pythagorean Theorem

OpenStudy (triciaal):